# The Euclidean distance degree of curves: from rational to line multiview varieties

### v1

@@ Abstract
The Euclidean distance (ED) degree is an invariant that measures the algebraic complexity of optimizing the distance function of a point to a model. It has been studied in algebraic statistics, machine learning, and computer vision. In this article, we prove a formula for the ED degree of curves parameterized by rational functions with mild genericity assumptions. We apply our results to resolve conjectures on one-dimensional line multiview varieties from computer vision proposed by Duff and Rydell.
@@ Introduction
One of the most impactful applications of algebraic geometry is in computer vision. This has led to the development of what is now known as algebraic vision. A featured problem in this area is to determine the number of critical points of the triangulation and, more generally, the multiview triangulation problem in multiple view geometry [13]. This amounts to minimizing the Euclidean distance function to an algebraic set known as the (affine) multiview variety. Recent work by Duff and Rydell introduced an extensive catalogue of multiview varieties motivated by new vision applications. In turn, this led to questions and conjectures on the Euclidean distance degrees for these models. In this article, we prove two of these conjectures using a new result on the ED degree of curves in products of projective spaces.
@@
Our paper is structured as follows. In Section 1 we provide the necessary preliminaries for the tools we use to prove our main results. We also highlight prior work in the algebraic vision community to put our results into context and address some of the subtleties in the proofs. In Section 2 we prove a general statement on the Euclidean distance degree of the multiview variety of rational curves of arbitrary degree. In Section 3 we have the main application of our result: strengthening the catalogue theorem [7, Theorem 3.3] of Duff-Rydell by settling their conjectures on one-dimensional multiview varieties. In Section 4 we conclude with a brief discussion on future directions.
@@ 1.1 Multiprojective varieties and multidegrees
A subvariety of a product of projective spaces is known as a multiprojective variety. Let $Z\subset(\mathbb{P}^{h})^{n}$ be such a variety. Its defining ideal lies in the multigraded ring
@@
$\operatorname{Deg}(Z):=\sum_{\alpha_{1}+\cdots+\alpha_{n}=\dim{Z}}c_{\alpha}\frac{T_{1}^{h}\cdots T_{n}^{h}}{T^{\alpha}}$
in the quotient ring $\mathbb{Z}[T_{1},\dots,T_{n}]/\langle T_{1}^{h+1},\dots,T_{n}^{h+1}\rangle$. We use the standard multi-index notation here, so that if $\alpha=(\alpha_{1},\dots,\alpha_{n})$ then $T^{\alpha}:=T_{1}^{\alpha_{1}}\cdots T_{n}^{\alpha_{n}}$. The coefficient $c_{\alpha}$ is the number of intersection points of $Z$ with a general linear space of the form
@@
where $\ell_{i,j}$ is a general linear form in the unknowns $\overline{\mathbf{x}}_{i}$.
@@
Consider the bi-homogeneous polynomials
@@
and the multidegree of $V(f_{1},f_{2},f_{3})\cap V(f_{0})$ is $6T_{1}^{3}T_{2}+11T_{1}^{2}T_{2}^{2}+6T_{1}T_{2}^{3}.$ $\diamond$
@@ 1.2 Multiprojective versus affine multiview varieties
In multiview geometry, the natural setting is multiprojective space. However, optimization problems such as Euclidean distance minimization are typically formulated in an affine chart. We now describe both perspectives to get a multiprojective and an affine multiview variety.
@@
As in Section 1.1, we take the coordinates of $(\mathbb{P}^{2})^{n}$ to be $\overline{\mathbf{x}}_{1},\dots,\overline{\mathbf{x}}_{n}$ The multidegree of the image of (1.1) is
$\sum_{\max\{\alpha_{1},\dots,\alpha_{n}\}=2,\,\,\alpha_{1}+\cdots+\alpha_{n}=3}1\cdot{T_{1}^{2-\alpha_{1}}\cdot T_{2}^{2-\alpha_{2}}\cdot\cdots\cdot T_{n}^{2-\alpha_{n}}},\quad n>1.$ (1.2)
For a textbook introduction to multidegrees see [23, Chapter 8]. Ideals of multiview varieties are also well studied in algebraic vision, and for more details we highlight [1, 3, 9].
@@
In general, if $Z$ is a multiprojective variety then we denote its restriction to the conventional affine chart by $Z_{{\operatorname{Aff}}}\subset(\mathbb{K}^{h})^{n}$.
@@ 1.3 Euclidean distance degree
The Euclidean distance degree (ED degree) of an algebraic variety measures the algebraic complexity of the problem of minimizing Euclidean distance to that variety. This invariant plays a central role in applications such as computer vision, statistics, and optimization. We recall its definition in this subsection.
@@
has first-order optimality conditions. Namely, for a smooth point $x\in X_{\mathbb{R}}$, we require $(x-u)^{T}v=0$ for all $v$ in the tangent space $T_{x}X_{\mathbb{R}}$ of $X_{\mathbb{R}}$ at $x$. This condition gives rise to a polynomial system [6, Equation 2.1] that has $x$ as the unknown and $u$ as the given generic data. The number of complex (real or nonreal) solutions to this polynomial system is called the Euclidean distance degree of $X_{\mathbb{R}}$ [6]. The ED degree is a measure of complexity for minimizing the Euclidean distance from $X_{\mathbb{R}}$ to $u$.
@@
where (a,b,c,s)=(0.8,0.74,0.5,0.32)(a,b,c,s)=(-0.8,-0.74,-0.5,0.32). Then, for a generic data point $u$, the ideal of the critical points is generated by $g_{1},g_{2}$ and the determinant of
@@
Figure 2: The red curve is $X_{\mathbb{R}}$ and the real ED critical points for the distance function to the orange points is the set of intersection points with the silver curve.
@@
For many of our purposes, $X$ is smooth and it suffices to consider the Euler-Poincare characteristic $\chi$.
@@
In Section 2 we will apply these results to compute the ED degree of one dimensional multiview varieties.
@@ 1.4 The projective ED degree
The projective ED degree provides a natural way to study ED degrees of affine cones by avoiding the singularity at the cone point and enabling the use of intersection-theoretic tools. In this subsection, we recall this invariant and prior work for curves.
Let $Z\subset\mathbb{P}^{h}$ be an irreducible closed subvariety of codimension $c$ and let $I_{Z}$ be its defining ideal. If the homogeneous polynomials $g_{1},\dots,g_{s}\in\mathbb{C}[z_{0},z_{1},\dots,z_{h}]$ generate $I_{Z}$, then the projective Euclidean distance degree correspondence is a biprojective variety in $\mathbb{P}^{h}\times\mathbb{P}^{h}$ defined by the saturated ideal
@@
The ED bidegree is defined to be multidegree of the ED degree correspondence. The coefficient of $T_{2}^{h}$ is the (projective) ED degree of $X$, and is denoted by $\operatorname{EDdeg_{proj}}(X)$. The other coefficients of the multidegree are known as the sectional ED degree, and the degree of $X$ is the coefficient of $T_{1}^{c}T_{2}^{h-c}$.
@@
Theorem 1.5 (Theorem 1.3 [21]).
Let $Z$ be any irreducible closed subvariety of $\mathbb{P}^{h}$. Then for a general ${{\boldsymbol{\beta}}}=(\beta_{0},\ldots,\beta_{h})\in{{\mathbb{C}}}^{{h+1}}\setminus\{0\}$ we have
@@
Consider the curve $Z$ in $\mathbb{P}^{2}\times\mathbb{P}^{2}$ parameterized by
$\mathbb{P}^{1}\to Z\subset\mathbb{P}^{2}\times\mathbb{P}^{2}$
@@
$\frac{9}{2}n^{3}-\frac{21}{2}n^{2}+8n-4.$
However, by [12, Theorem 1], its projective closure in $\mathbb{P}^{2n}$ has ED degree
@@ 1.5 More multiview varieties and triangulation
In this subsection, we define what are known as anchored point multiview varieties. To stay organized and minimize notation, we focus on the camera arrangement and the image it induces.
@@
where the overline denotes the Zariski closure. If $Y\subseteq\mathbb{P}^{N}$ is a subvariety, we define
@@
When $N=3$ and $h=2$, the map $\Phi_{\mathbf{C}}$ models the procedure of taking $n$ pictures with $n$ possibly distinct cameras. A point $P\in\mathbb{P}^{3}$ is mapped to the tuple of images $(C_{1}P,\dots,C_{n}P)$ provided that $P$ is not in the kernel of any $C_{i}$. Points $P\in\mathbb{P}^{N}$ that lie in the kernel of some camera are referred to as being at a camera center, where $\Phi_{\mathbf{C}}$ is undefined. That is, the camera centers constitute the base locus of $\Phi_{\mathbf{C}}$.
Example 1.10 (A one dimensional multiview variety).
If $Y$ is a line in $\mathbb{P}^{3}$, consider the anchored multiview variety $(C_{1},\dots,C_{n})\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y$. The affine multiview variety of $(C_{1},\dots,C_{n})\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y$ has ED Degree 3n23n-2. This result appears in [24, Theorem 1.7] and serves as a baseline case for understanding the complexity of triangulating points on curves. $\diamond$
@@
where $X_{\operatorname{Aff}}$ is the restriction of $X$ to the conventional affine chart described in Section 1.2.
@@ 1.6 A Schubert variety in $\operatorname{Gr}(1,\mathbb{P}^{3})$
In this subsection, we define a subvariety of the Grassmannian that will induce the multiprojective varieties appearing in our results in Section 3. Let $\operatorname{Gr}(1,\mathbb{P}^{3})$ denote the Grassmannian of lines in $\mathbb{P}^{3}$. Given three skew lines ${\ell^{(1)}},{\ell^{(2)}},{\ell^{(3)}}$ we define
@@
$\iota:\operatorname{Gr}(1,\mathbb{P}^{3})\to\mathbb{P}^{5},$
@@
with a linear map $B_{\texttt{L}^{\texttt{3}}}$ represented by the matrix
$B_{{\texttt{L}^{\texttt{3}}}}=\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&-1&0\\ 0&1&0\\ 0&0&0\\ 0&0&1\end{bmatrix}$
so that the parameterization map $f:\mathbb{P}^{1}\to\iota({\texttt{L}^{\texttt{3}}})$ takes $[s:t]$ to $B_{{\texttt{L}^{\texttt{3}}}}\cdot\nu([s:t])$.
@@
If $\mathbf{D}$ is a camera arrangement of size $(H+1)\times 6$, then $\mathbf{D}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\iota({\texttt{L}^{\texttt{3}}})\subset(\mathbb{P}^{H})^{n}$ is a point multiview variety anchored at $\iota({\texttt{L}^{\texttt{3}}})$. This variety was previously studied in [7] when $\mathbf{D}$ is an arrangement of so-called wedge cameras (Section 3.2). We revisit this construction and show how it relates to line multiview varieties in Section 3 for $H=\binom{2+1}{2}-1$ and $H=\binom{3+1}{2}-1$.
2 ED degree of multiview varieties of a rational curve

### v2

# The Euclidean distance degree of one-parameter anchored multiview varieties

@@ Abstract
Multiview varieties are mathematical models for the set of image feature correspondences that can be produced by a given camera arrangement. They possess an invariant known as their Euclidean distance (ED) degree, which measures the algebraic complexity of determining the 3D features that minimize the reprojection error when computing the scene structure by triangulation. In this article, we prove a formula for the ED degree of curves parameterized by rational functions with mild genericity assumptions. We apply our results to resolve conjectures on one-dimensional line multiview varieties from computer vision proposed by Duff and Rydell.
@@ Introduction
One of the most impactful applications of algebraic geometry is in computer vision, leading to the development of the field now known as algebraic vision. A featured problem in this area is to determine the number of critical points arising in triangulation and, more generally, in multiview triangulation in multiple view geometry (see [19] for a textbook reference). This amounts to minimizing a least squares reprojection error, and is mathematically framed as minimizing the squared Euclidean distance function to an algebraic set known as the (affine) multiview variety. Recent work by Duff and Rydell introduced an extensive catalogue of multiview varieties motivated by reconstruction applications [12, 18, 35, 31] and recent theoretical advances [4, 5, 20, 32, 33]. In turn, this led to questions and conjectures on the Euclidean distance degrees for these models. In this article, we prove two of these conjectures using a new result on the ED degree of curves in products of projective spaces.
@@
Our paper is structured as follows. In Section 1 we provide the necessary preliminaries for the tools we use to prove our main results. We also highlight prior work in the algebraic vision community to put our results into context and address some of the subtleties in the proofs. In Section 2 we prove a general statement on the Euclidean distance degree of the multiview variety of a rational curve of arbitrary degree. In Section 3 we have the main application of our result: settling the conjectures [9, Conjecture 7.4.5 and Conjecture 7.4.6] of Duff–Rydell on one-dimensional multiview varieties. In Section 4 we apply Section 2 to determine the ED degree of a class of one-parameter families of lines in the Grassmannian. In Section 5 we conclude with a brief discussion on future directions.
@@ 1.1 Multiprojective varieties and multidegrees
Solution sets of polynomial equations are algebraic varieties. When each polynomial is homogeneous, the solution set is invariant under rescaling, which gives rise to a projective variety. When the defining equations are homogeneous with respect to multiple independent rescalings, we obtain a subvariety of a product of projective spaces known as a multiprojective variety. In this subsection we review some basics of multiprojective varieties and their multidegrees, and in the next section we see how they arose in computer vision as multiview varieties.
Let $Z\subset(\mathbb{P}^{h})^{n}$ be such a variety. Its defining ideal lies in the multigraded ring
@@
$\operatorname{Deg}(Z):=\sum_{\alpha_{1}+\cdots+\alpha_{n}=\dim{Z}}c_{\alpha}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{T_{1}^{h-\alpha_{1}}\cdots T_{n}^{h-\alpha_{n}}}}$
in the quotient ring $\mathbb{Z}[T_{1},\dots,T_{n}]/\langle T_{1}^{h+1},\dots,T_{n}^{h+1}\rangle$. This quotient reflects the geometric fact that one cannot have a nonempty intersection of a factor $\mathbb{P}^{h}$ with more than $h$ general hyperplanes. We will use the standard multi-index notation, so that if $\alpha=(\alpha_{1},\dots,\alpha_{n})$ then $T^{\alpha}:=T_{1}^{\alpha_{1}}\cdots T_{n}^{\alpha_{n}}$. The coefficient $c_{\alpha}$ is the number of intersection points of $Z$ with a general linear space of the form
@@
where $\ell_{i,j}$ is a general linear form in the unknowns $\overline{\mathbf{x}}_{i}$. One should think of the exponents of the terms of the polynomial as recording how many hyperplanes we impose in each factor, and of the coefficients as counting the size of the corresponding intersection.
@@
The following shows how multidegrees behave under intersecting bihomogeneous hypersurfaces. Consider the bi-homogeneous polynomials
@@
and the multidegree of $V(f_{1},f_{2},f_{3})\cap V(f_{0})$ is $6T_{1}^{3}T_{2}+11T_{1}^{2}T_{2}^{2}+6T_{1}T_{2}^{3}.$
This example shows how multidegrees behave under intersecting bihomogeneous hypersurfaces. $\diamond$
@@ 1.2 Multiprojective versus affine multiview varieties
In multiview geometry, the natural setting is multiprojective space. However, optimization problems such as Euclidean distance minimization are typically formulated in an affine chart. We now describe how multiview varieties appear in both multiprojective and affine settings. In computer vision, these encode the algebraic constraints satisfied by feature correspondences across multiple images.
@@
As in Section 1.1, we take the coordinates of $(\mathbb{P}^{2})^{n}$ to be $\overline{\mathbf{x}}_{1},\dots,\overline{\mathbf{x}}_{n}$. The multidegree of the image of (1.1) is
$\sum_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\max\{\alpha_{1},\dots,\alpha_{n}\}\leq 2,}\,\,\alpha_{1}+\cdots+\alpha_{n}=3}1\cdot{T_{1}^{2-\alpha_{1}}\cdot T_{2}^{2-\alpha_{2}}\cdot\cdots\cdot T_{n}^{2-\alpha_{n}}},\quad n>1.$ (1.2)
The exponent $2-\alpha_{i}$ indicates that $\alpha_{i}$ hyperplanes were used in the $i$th factor. Ideals and multidegrees of multiview varieties are also well studied in algebraic vision, and for more details we highlight [1, 5, 11]. For a textbook introduction to multidegrees, see [30, Chapter 8].
@@
In general, if $Z$ is a multiprojective subvariety in $(\mathbb{P}^{h})^{n}$ then we denote its restriction to the conventional affine chart by $Z_{{\operatorname{Aff}}}\subset(\mathbb{K}^{h})^{n}$. This corresponds to working with inhomogeneous image coordinates, where reprojection error is naturally defined.
@@ 1.3 Euclidean distance degree
The Euclidean distance degree (ED degree) of an algebraic variety measures the algebraic complexity of the problem of minimizing Euclidean distance to that variety. We recall its definition in this subsection.
@@
has first-order optimality conditions. Namely, for a smooth point $x\in X_{\mathbb{R}}$, we require $(x-u)^{T}v=0$ for all $v$ in the tangent space $T_{x}X_{\mathbb{R}}$ of $X_{\mathbb{R}}$ at $x$. This condition gives rise to a polynomial system [8, Equation 2.1] that has $x$ as the unknown and $u$ as the given generic data. The number of complex (real or nonreal) solutions to this polynomial system is called the Euclidean distance degree of $X_{\mathbb{R}}$ [8]. The ED degree is a measure of complexity for minimizing the squared Euclidean distance from $X_{\mathbb{R}}$ to $u$.
@@
where (a,b,c,s)=(0.8,0.74,0.5,0.32)(a,b,c,s)=(-0.8,-0.74,-0.5,0.32). Then, for a generic data point $u$, the ideal of the critical points is generated by $g_{1},g_{2}$ and the determinant of the $3\times 3$ matrix
@@
Figure 2: The red curve is $X_{\mathbb{R}}$ and the real ED critical points for the distance function to the orange point is the set of intersection points with the silver curve.
@@
For many of our purposes, $X$ is smooth and it suffices to consider the Euler–Poincaré characteristic $\chi$.
@@
We apply these results to compute the ED degrees of one-dimensional multiview varieties in Section 2 and resolve computer vision conjectures in Section 3.
@@ 1.4 The projective ED degree
The projective ED degree provides a natural way to study ED degrees of affine cones by avoiding the singularity at the cone point and enabling the use of intersection-theoretic tools. In this subsection, we recall this invariant and prior work for curves. Although our main results concern affine ED degrees, the projective ED degree also plays an important role because many multiview varieties arise as affine cones or have natural projective closures.
Let $Z\subset\mathbb{P}^{h}$ be an irreducible closed subvariety of codimension $c$ and let $I_{Z}$ be its defining ideal. If the homogeneous polynomials $g_{1},\dots,g_{s}\in\mathbb{C}[z_{0},z_{1},\dots,z_{h}]$ generate $I_{Z}$, then the projective Euclidean distance correspondence is a biprojective variety in $\mathbb{P}^{h}\times\mathbb{P}^{h}$ defined by the saturated ideal
@@
The ED bidegree is defined to be multidegree of the ED correspondence. The coefficient of $T_{2}^{h}$ is the (projective) ED degree of $X$, and is denoted by $\operatorname{EDdeg_{proj}}(X)$.
@@
Theorem 1.5.
[28, Theorem 1.3] Let $Z$ be any irreducible closed subvariety of $\mathbb{P}^{h}$. Then for a general ${{\boldsymbol{\beta}}}=(\beta_{0},\ldots,\beta_{h})\in{{\mathbb{C}}}^{{h+1}}\setminus\{0\}$ we have
@@
Consider the curve $Z$ in $\mathbb{P}^{2}\times\mathbb{P}^{2}$ parameterized by the map
$\mathbb{P}^{1}\to Z\subset\mathbb{P}^{2}\times\mathbb{P}^{2},$
@@
The ED degree is not invariant under projective closure. Even for classical multiview varieties, the affine and projective versions may have drastically different ED degrees.
@@
$\frac{9}{2}n^{3}-\frac{21}{2}n^{2}+8n-4.$ (1.5)
This cubic polynomial in $n$ counts the number of (real or nonreal) critical points of the distance function from a general data point to an $n$-view variety when the cameras are in generic position. The values of this polynomial for small $n$ was determined by [34]. In particular, for $n=3$, there are forty seven critical points, and [34] showed how to minimize the distance function to a three-view variety by solving a $47\times 47$ eigenvalue problem.
In [8], it was noticed that the projective closure in $\mathbb{P}^{5}$ of a three-view variety has ED degree 56. In [17, Theorem 1], they found the projective closure of the $n$-view variety in $\mathbb{P}^{2n}$ has ED degree
@@ 1.5 More multiview varieties and triangulation
Anchored multiview varieties arise naturally when the world contains known structures, such as a curve, line, or surface. These models describe all possible image observations of points constrained to lie on $Y$. In this subsection, we recall the definition of anchored point multiview varieties. To stay organized and minimize notation, we focus on the camera arrangement and the image it induces.
@@
where the overline denotes the Zariski closure. If $Y\subseteq\mathbb{P}^{N}$ is a subvariety, define
@@
When $N=3$ and $h=2$, the map $\Phi_{\mathbf{C}}$ models the procedure of taking $n$ pictures with $n$ possibly distinct cameras. A world point $P\in\mathbb{P}^{3}$ is mapped to the tuple of images $(C_{1}P,\dots,C_{n}P)$ provided that $P$ is not in the kernel of any $C_{i}$. Points $P\in\mathbb{P}^{N}$ that lie in the kernel of some camera are referred to as being at a camera center, where $\Phi_{\mathbf{C}}$ is undefined. That is, the camera centers constitute the base locus of $\Phi_{\mathbf{C}}$.
Example 1.10 (A one-dimensional multiview variety).
If $Y$ is a line in $\mathbb{P}^{3}$, consider the anchored multiview variety $(C_{1},\dots,C_{n})\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y$. The affine multiview variety of $(C_{1},\dots,C_{n})\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y$ has ED degree 3n23n-2. This result appears in [33, Theorem 1.7] and serves as a baseline case for understanding the complexity of triangulating points on curves. This is the degree one case of Theorem 2.3. $\diamond$
@@
where $X_{\operatorname{Aff}}$ is the restriction of $X$ to the affine chart described in Section 1.2.
@@ 1.6 A Schubert variety in $\operatorname{Gr}(1,\mathbb{P}^{3})$
In this subsection, we consider a subvariety of the Grassmannian that will induce the multiprojective varieties appearing in our results in Section 3. Let $\operatorname{Gr}(1,\mathbb{P}^{3})$ denote the Grassmannian of lines in $\mathbb{P}^{3}$. Given three skew lines ${\ell^{(1)}},{\ell^{(2)}},{\ell^{(3)}}$ we have
@@
$\iota:\operatorname{Gr}(1,\mathbb{P}^{3})\to\mathbb{P}^{5}$
@@
with a linear map, denoted by $B_{\texttt{L}^{\texttt{3}}}$, so that the parameterization map $f:\mathbb{P}^{1}\to\iota({\texttt{L}^{\texttt{3}}})$ takes $[s:t]$ to $B_{{\texttt{L}^{\texttt{3}}}}\cdot\nu([s:t])$.
@@
If $\mathbf{D}$ is an arrangement of cameras of size $(H+1)\times 6$, then $\mathbf{D}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\iota({\texttt{L}^{\texttt{3}}})\subset(\mathbb{P}^{H})^{n}$ is a point multiview variety anchored at $\iota({\texttt{L}^{\texttt{3}}})$. This variety was previously studied in [9] when $\mathbf{D}$ is an arrangement of so-called wedge cameras (Section 3.2). We revisit this construction and show how it relates to line multiview varieties in Section 3 for $H=\binom{2+1}{2}-1$ and $H=\binom{3+1}{2}-1$.
2 ED degrees of multiview varieties anchored at rational curves